OpenStudy (anonymous):

Jason has already read 27 pages of a book. He plans to read 3 pages of the book every day. If f(n) represents the total number of pages that Jason will complete reading in n days, which of the following functions represents the relationship between f(n) and n? A. f(n) = 27n - 3 B. f(n) = 3n - 27 C. f(n) = 27n + 3 D. f(n) = 3n + 27

OpenStudy (anonymous):

PLZ i need help

jimthompson5910 (jim_thompson5910):

Day 0: 27 pages (this is what we start with) Day 1: 27+3 = 30 pages (end of day 1, add 3 pages) Day 2: 27+3+3 = 33 pages (end of day 2, add 3 more pages) Day 3: 27+3+3+3 = 36 pages (end of day 3, add 3 more pages) Day 4: 27+3+3+3+3 = 39 pages (end of day 3, add 3 more pages) etc etc

jimthompson5910 (jim_thompson5910):

So how many pages are read on Day n?

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

that's the number of pages added per day I'm asking how many pages total are read on day n?

OpenStudy (anonymous):

what it day n

jimthompson5910 (jim_thompson5910):

How many times did we add 3 for day 1?

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

How many times did we add 3 for day 2?

OpenStudy (anonymous):

would the answer be C

jimthompson5910 (jim_thompson5910):

How many times did we add 3 for day 2?

OpenStudy (anonymous):

on day 2 we added 2 and on day 3 we added 3 and on day 4 we added 4

jimthompson5910 (jim_thompson5910):

Good. So on any day, call it day x, we add 3 a total of x times 3+3+3+....+3 <--- Adding up 3...x number of times we shorten this up to 3x (3 times x) Then we add on 27, which is the base number we already started with to get 3x + 27

jimthompson5910 (jim_thompson5910):

Now because they're using n in place of x, we just replace x with n to get this final answer f(n) = 3n + 27

jimthompson5910 (jim_thompson5910):

As confirmation, you can plug in say n = 2 (for day 2) to get f(n) = 3n + 27 f(2) = 3*2 + 27 f(2) = 6 + 27 f(2) = 33 And it matches up with the info above

OpenStudy (anonymous):

so the answer would be D

jimthompson5910 (jim_thompson5910):

Correct

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome